Show that the work necessary to assemble a static charge distribution (against the coulomb repulsion of like charges)
1
Expert's answer
2021-10-25T00:15:01-0400

Assume charge is static (q)

Force repulsion mean

"q_1=q,q_2=q"

Between charge distance =r

"F=\frac{kq_1q_2}{r^2}"

Put value

"F=\frac{kq^2}{r^2}" Repulsion force

Work (W)

"W=-\frac{kq_1q_2}{r}"

"W=-\frac{kq^2}{r}"

∆U=W="\frac{kq^2}{r}"